OpenStudy (anonymous):

will medal need help The ratio of men to women working for a company is 7 to 5 . If there are 264 employees total, how many women work for the company?

OpenStudy (freckles):

you are given that the sum of #men and #women is 264. you are given that #men/#women=7/5 Let's use m for the number of men And w for the number of women. So you are given: \[m+w=264 \\ \frac{m}{w}=\frac{7}{5} \implies m=\frac{7}{5}w \text{ plug this into first equation }\]

OpenStudy (anonymous):

110